# Crypto Inverse-Power Options and Fractional Stochastic Volatility

### v2

@@ 6 Conclusions
and, with $\varDelta V_{S}(s,T)=H(T,s)\varDelta Y_{s}\mathsf{Var}[X_{1}]$ from (2.2.2) and (Appendix A),
$\varDelta\phi_{\log S_{T}|\cdot})_{s}(u)=\phi_{\log S_{T}|s-}(u)\big{(}\exp\big{(}{\rm i}u(\varDelta X_{\mathcal{T}_{s}}+\rho\varDelta Y_{s})+(\log\phi_{X_{1}}(u)-{\rm i}u\phi_{X_{1}}(-{\rm i}))H(T,s)\varDelta Y_{s}\big{)}-1\big{)}.$ (A.8)
By substituting (Appendix A) and (A.8) into (3.1.9), all the items in (1) follow immediately.
It remains to justify for (3.1.9) the validity of differentiation under the integral sign, while passing it inside the real part is allowed by the Hermitian property of characteristic functions. Indeed, since $|K^{-{\rm i}up_{2}/p_{1}}|=1$ and $\int^{\infty}_{0}\mathrm{Re}[\phi_{\log S_{T}|s}(u)]{\rm d}u$ converges by the absolute continuity of $\mathcal{L}(\log S_{T}|\mathscr{F}_{s})$ (Theorem 1), we only need to consider the tail integrability of the multiplier of $K^{-{\rm i}up_{2}/p_{1}}\phi_{\log S_{T}|s}(u)$ in each item in (1).
@@
from (4.2.4). To compute the integral in (1) involving $\log\phi_{Y_{1}}$, it suffices to focus on the integral of $(b_{Y}-{\rm i}(\rho u+H_{3}(t-s)\varphi(u)))^{c_{Y}}$, where $\varphi$ is as defined in (4.2). To this end, by letting $H_{3,-}$ and $H_{3,+}$ be the first piece and second piece, respectively, of $H$ in (4.1.3), direct integration yields
@@
For the first two integrals in the last equality of (Appendix A), take a placeholder $\spadesuit\in\{t-t_{0},(1-d)/\kappa\}\subsetneq\mathbb{R}_{++}$; the substitution $s^{d}\mapsto s$ then gives
@@
$=-\frac{\varphi^{c_{Y}}_{3}(u)}{\kappa}\bigg{(}\frac{\varphi_{3}(u)/(s\varphi_{4}(u))-1}{1-\varphi_{3}(u)/(s\varphi_{4}(u))}\bigg{)}^{c_{Y}}\mathrm{B}_{\varphi_{3}(u)/(s\varphi_{4}(u))}(-c_{Y},c_{Y}+1)\bigg{|}^{e^{-(1-d)}}_{s=e^{-\kappa(t-t_{0})}}$
@@
For the case $c_{Y}=0$, it is quite cumbersome to derive the results as direct limits of the above; instead, we perform a separate computation. With $c_{Y}=0$ and the procedures in (Appendix A) in mind, we are led to evaluate the integrals of $\log(\varphi_{1}(u)+s^{d}\varphi_{2}(u))$ and $\log(\varphi_{3}(u)-e^{-\kappa s}\varphi_{4}(u))$, respectively. First, by expanding the logarithm around $s=0$, we have

### v3

@@ Appendix A Mathematical proofs
and, with $\varDelta V_{S}(s,T)=H(T,s)\varDelta Y_{s}\mathsf{Var}[X_{1}]$ from (2.2.2) and (A),
$(\varDelta\phi_{\log S_{T}|\cdot})_{s}(u)=\phi_{\log S_{T}|s-}(u)\big{(}\exp\big{(}{\rm i}u(\varDelta X_{\mathcal{T}_{s}}+\rho\varDelta Y_{s})+(\log\phi_{X_{1}}(u)-{\rm i}u\phi_{X_{1}}(-{\rm i}))H(T,s)\varDelta Y_{s}\big{)}-1\big{)}.$ (A.8)
By substituting (A) and (A.8) into (3.1.9), all the items in (1) follow immediately.
It remains to justify for (3.1.9) the validity of differentiation under the integral sign, while passing it inside the real part is allowed by the Hermitian property of characteristic functions. Indeed, since $|K^{-{\rm i}up_{2}/p_{1}}|=1$ and $\int^{\infty}_{0}\mathrm{Re}[\phi_{\log S_{T}|s}(u)]{\rm d}u$ converges by the absolute continuity of $\mathcal{L}(\log S_{T}|\mathcal{F}_{s})$ (Theorem 1), we only need to consider the tail integrability of the multiplier of $K^{-{\rm i}up_{2}/p_{1}}\phi_{\log S_{T}|s}(u)$ in each item in (1).
@@
from (4.2.4). To compute the integral in (1) involving $\log\phi_{Y_{1}}$, it suffices to focus on the integral of $(b_{Y}-{\rm i}(\rho u+H_{3}(t-s)\varphi(u)))^{c_{Y}}$, where $\varphi$ is as defined in (4.2). By letting $H_{3,-}$ and $H_{3,+}$ be the first piece and second piece, respectively, of $H$ in (4.1.3), direct integration yields
@@
For the first two integrals in the last equality of (A), take a placeholder $\spadesuit\in\{t-t_{0},(1-d)/\kappa\}\subsetneq\mathbb{R}_{++}$; the substitution $s^{d}\mapsto s$ then gives
@@
$=-\frac{\varphi^{c_{Y}}_{3}(u)}{\kappa}\bigg{(}\frac{\varphi_{3}(u)/(s\varphi_{4}(u))-1}{1-\varphi_{3}(u)/(s\varphi_{4}(u))}\bigg{)}^{c_{Y}}$
×$\qquad\times\mathrm{B}_{\varphi_{3}(u)/(s\varphi_{4}(u))}(-c_{Y},c_{Y}+1)\bigg{|}^{e^{-(1-d)}}_{s=e^{-\kappa(t-t_{0})}}$
@@
For the case $c_{Y}=0$, it is quite cumbersome to derive the results as direct limits of the above; instead, we perform a separate computation. With $c_{Y}=0$ and the procedures in (A) in mind, we are led to evaluate the integrals of $\log(\varphi_{1}(u)+s^{d}\varphi_{2}(u))$ and $\log(\varphi_{3}(u)-e^{-\kappa s}\varphi_{4}(u))$, respectively. First, by expanding the logarithm around $s=0$, we have